OpenStudy (anonymous):

the slope of a line is -1\2. the point (6,-7) is on the line. Determine the equation of this line.

OpenStudy (anonymous):

slope(m) = -1/2 (6,-7) x = 6 and y = -7 y = mx + b (solving for b) now we sub y = mx + b -7 = -1/2(6) + b -7 = -3 + b -7 + 3 = b -4 = b your equation is : y = -1/2x - 4

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem :)